OpenStudy (anonymous):

what are the values of x when x+2sincos=1?

OpenStudy (michele_laino):

please, I don't understand your equation, please can you rewrite it?

OpenStudy (anonymous):

well the original equation was x+sin(x)^2 and I have to find where f prime is equal to one

OpenStudy (michele_laino):

Ok!, I write: \[f(x)=x+(\sin x)^{2}\] then \[f'(x)=1+2 \sin x \cos x=1+\sin(2x)\] from the condition: \[f'(x)=1\] we get: \[1+\sin(2x)=1\] so: \[\sin(2x)=0\] finally: \[2x=k \pi\] or: \[x=\frac{ k \pi }{ 2 }\] where: \[k=0,\pm1,\pm2,\pm3,...\]

OpenStudy (anonymous):

where did the kpi come from?

OpenStudy (michele_laino):

generally the solution of equation: \[\sin \alpha=0\] are: \[\alpha = k \pi\] because in the unit circle, the points which have y-coordinate equal to zero are the only point located at 0°, 180°, 360°, and so on, whereas if we go in the negative sense, we have 0°, -180°, -360°, and so on